OpenStudy (anonymous):

-150=290.8-190e^(-.75x) Can any one help me solve this equation please? @precal

OpenStudy (tkhunny):

Are we to solve for x? Subtract 290.8 Divide by -190 Introduce Logarithms. What else?

OpenStudy (anonymous):

Ok .. I got it down to 2.32=e^(-.75x) I am not for sure what to do next. @tkhunny

OpenStudy (kropot72):

\[-150=290.8-190e ^{-0.75x}\] \[190e ^{-0.75x}=290.8+150=440.8\] \[e ^{-0.75x}=\frac{440.8}{190}\] Taking natural logs of both sides gives: \[-0.75x=\ln \frac{440.8}{190}=0.841567185\] \[x=\frac{0.841567185}{-0.75}=-1.122\]

OpenStudy (precal):